# Perfect Wave Transfer in Continuous Quantum Systems

### v1

@@ 2. Example 2:
It is curious to note the following “shortcut” to writing our massless bosons (8) in Liouville normal form: Recall that an inhomogeneous TLL is a theory of compactified free bosons with a position-dependent compactification radius $R(x)$ in curved spacetime given by the metric $(h_{\mu\nu})=\operatorname{diag}(v(x)^{2}/v_{0}^{2},-1)$ for coordinates $(x^{0},x^{1})\equiv(v_{0}t,x)$; see, e.g., [27, 53]. Given a Luttinger parameter profile $K(x)$, the compactification radius is $R(x)=\sqrt{2\alpha^{\prime}}K(x)$ ($\alpha^{\prime}$ has dimension length squared and is commonly set equal to $2$). In terms of the bosonic field $\varphi(x)$ (with values taken modulo $2\pi$), the action corresponding to the Hamiltonian (8) is
$S=\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}R(x)^{2}(\partial_{\mu}\varphi)(\partial^{\mu}\varphi),$ (50)
where we recall that $h\equiv\det(h_{\mu\nu})$. This can be viewed as a non-exactly marginal deformation of a theory with constant compactification radius; cf. [53]. The corresponding action for $X(x)\equiv R(x)\varphi(x)$ becomes
$S=\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}\bigl{[}(\partial_{\mu}X)(\partial^{\mu}X)-\hat{q}X^{2}/v(x)^{2}\bigr{]}$ (51)
with $\hat{q}=\hat{q}(y)$ in Eq. (18) for $y=y(x)$ given by Eq. (7), which agrees with Eq. (19) for $\mathfrak{u}(y)=X(x)/\sqrt{2\alpha^{\prime}}$. I.e., as a theory for $X(x)$, even when $q(x)=0$, the position-dependent $K(x)$ corresponds a (possibly imaginary) inhomogeneous mass $\sqrt{\hat{q}(y)}/v(x)$. This also extends to the case
$S=\frac{1}{2\pi}\int\mathrm{d}^{2}x\sqrt{-h}\bigl{[}K(x)(\partial_{\mu}\varphi)(\partial^{\mu}\varphi)-v_{0}^{2}M(x)^{2}\varphi^{2}\bigr{]}$ (52)
with a non-zero mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ for the bosonic field $\varphi(x)$, corresponding to the Hamiltonian (49), the only difference being that the first term in Eq. (18) is then no longer zero.
@@
Consider the inhomogeneous TLL theory with Hamiltonian $H_{\textrm{iTLL}}$ in Eq. (8). Using the eigenfunctions obtained from SL theory, we can compute any ground-state two-point correlation functions for the fields $\varphi(x)$ and $\vartheta(x)$, where the latter is defined by
@@
Indeed, using the expansions (45), that $a_{n}$ annihilates the ground state $|\Omega\rangle$ of $H_{\textrm{iTLL}}$, as well as the commutation relations $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$, it is straightforward to show that
@@
Similar correlation functions can be computed for other objects, including vertex operators (i.e., normal-ordered exponentials of the bosonic fields) [54, 55, 56, 57, 58, 59]. It is also straightforward to compute Green’s functions; see, e.g., [27], the only difference with the present case being that one must account for the non-trivial weight function $w(x)$ that was simply a constant in [27].
As an important example, consider a theory with $K(x)=K>0$ constant but $v(x)>0$ general [and $q(x)=0$]. We recall that the SL operator in this case is $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using Eq. (7) with eigenenergies $E_{n}=\pi nv_{0}/L$ and eigenfunctions $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl{(}n\pi[L+2y]/2L\bigr{)}$ and $u_{0}(x)=\sqrt{v_{0}/L}$. Then, one can show that
$\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle\\ \begin{aligned} =\frac{1}{4}&\left(\log\frac{\mathrm{i}\exp\Bigl{(}-\mathrm{i}\frac{\pi}{2L}[f(x)-f(x^{\prime})-v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)-f(x^{\prime})-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr{)}}+\log\frac{\mathrm{i}\exp\Bigl{(}-\mathrm{i}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L-v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr{)}}\right.\\ &\quad+\left.\log\frac{-\mathrm{i}\exp\Bigl{(}\mathrm{i}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L+v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr{)}}+\log\frac{-\mathrm{i}\exp\Bigl{(}\mathrm{i}\frac{\pi}{2L}[f(x)-f(x^{\prime})+v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)-f(x^{\prime})+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr{)}}\right),\end{aligned}$ (56)
where we included the necessary regularizations in order to interpret coincident spacetime points and used the identity $\sum_{n=1}^{\infty}\frac{1}{n}\mathrm{e}^{-\xi n}=-\log\left(1-\mathrm{e}^{-\xi}\right)$ for any $\xi$ such that $\operatorname{Re}\xi>0$. Eq. (56) is the analytical expression for what was plotted in Fig. 1 (assuming $K=1$). Moreover, for delta-like waves, the terms appearing in $\log F(t)$ given by Eq. (41) for the free-fermion case (37) closely resemble those in Eq. (56). Indeed, the former agrees with $4\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle$ to leading order, and one can furthermore show that the agreement between the two ways of computing (CFT methods versus bosonization) is exact (as it should be) by evaluating and summing all combinations of two-point correlation functions of the vertex operators describing the right/left-moving free fermionic fields.

### v3

@@ 2. Example 2:
It is curious to note the following “shortcut” to writing our massless boson model (8) in Liouville normal form: Recall that inhomogeneous TLL theory is that of compactified free bosons in curved spacetime with a position-dependent compactification radius $R(x)$ and the metric $(h_{\mu\nu})=\operatorname{diag}(v(x)^{2}/v_{0}^{2},-1)$ in coordinates $(x^{0},x^{1})\equiv(v_{0}t,x)$; see, e.g., [37, 61]. Given a Luttinger parameter profile $K(x)$, the compactification radius is $R(x)\equiv\sqrt{2\alpha^{\prime}}K(x)$ ($\alpha^{\prime}$ has dimensions of length squared and is commonly set equal to $2$). In terms of the bosonic field $\varphi(x)$ (with values taken modulo $2\pi$), the action corresponding to the Hamiltonian (8) is
$S_{\textrm{iTLL}}\equiv\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}R(x)^{2}(\partial_{\mu}\varphi)(\partial^{\mu}\varphi),$ (62)
where we recall that $h\equiv\det(h_{\mu\nu})$. This can be viewed as a nonexactly marginal deformation of a theory with constant compactification radius; cf. [61]. The corresponding action for $X(x)\equiv R(x)\varphi(x)$ becomes
$S_{\textrm{iTLL}}=\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}\bigl[(\partial_{\mu}X)(\partial^{\mu}X)-\hat{q}X^{2}/v(x)^{2}\bigr]$ (63)
with $\hat{q}=\hat{q}(y)$ in Eq. (18) for $y=y(x)$ given by Eq. (7), which agrees with Eq. (19) for $\mathfrak{u}(y)=X(x)/\sqrt{2\alpha^{\prime}}$. I.e., as a theory in terms of the field $X(x)$, even when $q(x)=0$, the position-dependent $K(x)$ corresponds a (possibly imaginary) inhomogeneous mass $\sqrt{\hat{q}(y)}/v(x)$. This also extends to the case
$S_{\textrm{miTLL}}\equiv\frac{1}{2\pi}\int\mathrm{d}^{2}x\sqrt{-h}\bigl[K(x)(\partial_{\mu}\varphi)(\partial^{\mu}\varphi)-v_{0}^{2}M(x)^{2}\varphi^{2}\bigr]$ (64)
with a nonzero mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ for the bosonic field $\varphi(x)$, corresponding to the Hamiltonian (61), the only difference being that the first term on the right-hand side of Eq. (18) is then no longer zero.
@@
Consider the inhomogeneous TLL theory with Hamiltonian $H_{\textrm{iTLL}}$ in Eq. (8). Using the eigenfunctions obtained from SL theory, we can compute any correlation function for the fields $\varphi(x)$ and $\vartheta(x)$, where the latter is defined by
@@
Case in point, using the expansions in Eq. (45) with zero modes omitted, that $a_{n}$ annihilates the ground state $|\Omega\rangle$ of $H_{\textrm{iTLL}}$, and the commutation relations $\bigl[a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr]=\delta_{n,n^{\prime}}$, it is straightforward to show that the ground-state two-point correlations are
@@
The first of these correlations is the same as the one in Eq. (17) in the main text. Similar correlations can be computed for other objects, including vertex operators, i.e., normal-ordered exponentials of the bosonic fields [62, 63, 64, 65, 66, 67]. It is also straightforward to compute Green’s functions; see, e.g., [37], the only difference with the present case being that one must account for the nontrivial weight $w(x)$ that was simply a constant in [37].
As an important example, consider a theory with $K(x)=K>0$ constant but $v(x)>0$ general [and $q(x)=0$]. We recall that the SL operator in this case is $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using $y=y(x)$ given by Eq. (7) with eigenenergies $E_{n}=\pi nv_{0}/L$ and eigenfunctions $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl(n\pi[L+2y]/2L\bigr)$ and $u_{0}(x)=\sqrt{v_{0}/L}$. One can then show that the expression (17) for the ground-state $\varphi\varphi$ correlation function (with zero modes omitted) yields
$\hskip-6.25958pt\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle\\ \begin{aligned} =\frac{1}{4}&\left(\log\frac{\mathrm{i}\exp\Bigl(-\mathrm{i}\frac{\pi}{2L}[y(x)-y(x^{\prime})-v_{0}(t-t^{\prime})]\Bigr)}{2\sin\Bigl(\frac{\pi}{2L}[y(x)-y(x^{\prime})-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr)}+\log\frac{\mathrm{i}\exp\Bigl(-\mathrm{i}\frac{\pi}{2L}[y(x)+y(x^{\prime})+L-v_{0}(t-t^{\prime})]\Bigr)}{2\sin\Bigl(\frac{\pi}{2L}[y(x)+y(x^{\prime})+L-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr)}\right.\\ &\quad+\left.\log\frac{-\mathrm{i}\exp\Bigl(\mathrm{i}\frac{\pi}{2L}[y(x)+y(x^{\prime})+L+v_{0}(t-t^{\prime})]\Bigr)}{2\sin\Bigl(\frac{\pi}{2L}[y(x)+y(x^{\prime})+L+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr)}+\log\frac{-\mathrm{i}\exp\Bigl(\mathrm{i}\frac{\pi}{2L}[y(x)-y(x^{\prime})+v_{0}(t-t^{\prime})]\Bigr)}{2\sin\Bigl(\frac{\pi}{2L}[y(x)-y(x^{\prime})+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr)}\right),\end{aligned}$ (68)
where we included necessary regularizations in order to interpret coincident spacetime points and used the identity $\sum_{n=1}^{\infty}\frac{1}{n}\mathrm{e}^{-\xi n}=-\log\left(1-\mathrm{e}^{-\xi}\right)$ for any $\xi$ such that $\operatorname{Re}\xi>0$. The real part of Eq. (68) gives the analytical expression for what was plotted in Fig. 1 (assuming $K=1$). Moreover, for deltalike waves, the terms appearing in $\log F(t)$ given by Eq. (41) for the free-fermion case (37) closely resemble those in Eq. (68). Indeed, the former agrees with $4\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle$ to leading order, and one can furthermore show that the agreement between the two ways of computing (CFT methods versus bosonization) is exact (as it should be) by evaluating and summing all combinations of two-point correlation functions of the vertex operators describing the right- and left-moving free fermionic fields.